# Models That Prove Their Own Correctness

### v3

@@ A.1 The general definition
Definition A.2 (Definition 3.2, generalized).
Fix a soundness error $s\in(0,1)$, a threshold $\lambda\in[0,1)$, a finite set of tokens $\Sigma$, and a loss function $\ell\colon\Sigma^{\ast}\times\Sigma^{\ast}\to[0,1]$. A verifier $V$ for $\ell$ with threshold $\lambda$ is a probabilistic polynomial-time algorithm that is given explicit inputs $x,y\in\Sigma^{\ast}$ and black-box (oracle) query access to a prover $P$. It interacts with $P$ over $R$ rounds (see Figure 1) and outputs a decision $V,P\right>(x,y)\in\{\mathrm{reject},\mathrm{accept}\}$. Verifier $V$ satisfies the following two guarantees:
@@
Indeed, for a given ground-truth function $F^{\ast}\colon\Sigma^{\ast}\to\Sigma^{\ast}$, Definition 3.2 can be recovered by choosing the 0-1 loss
$\ell_{F^{\ast}}(x,y)\coloneqq\begin{cases}1\quad\mathrm{if}\ x\neq F^{\ast}(y)\\ 0\quad\mathrm{else}.\end{cases}$
@@
Remark A.3 (Connection to Interactive Proofs of Proximity).
Definition A.2 can be seen as a slight generalization of (perfect completeness) Interactive Proofs of Proximity (IPPs, Rothblum et al. 2013). An IPP for a relation $L\subseteq\Sigma^{\ast}\times\Sigma^{\ast}$ with proximity parameter $\lambda$ is obtained by instantiating Definition A.2 with the loss function $\ell_{\mathrm{Hamming}}$ defined by
@@
that is, $\ell_{\mathrm{Hamming}}(x,y)$ is the fraction of tokens in $y$ that must be changed so as obtain an output $y^{\ast}$ with $(x,y^{\ast})\in L$. However, the motivation of Rothblum et al. [2013] was studying sublinear time verification, whereas ours is to relax the requirements of traditional Interactive Proofs towards meeting common desiderata in machine learning.
@@
Definition A.4 (Definition 3.4, generalized).
Fix a loss function $\ell\colon\Sigma^{\ast}\times\Sigma^{\ast}\to[0,1]$, a verifier $V$ for $\ell$ with threshold $\lambda\in[0,1)$ as in Definition A.2, and a distribution $\mu$ over inputs $\Sigma^{\ast}$. The Verifiability of a model $P_{\theta}\coloneq\Sigma^{\ast}\to\Sigma^{\ast}$ is defined as
@@
Analogously to Remark 3.5, we observe that Verifiability (Definition A.4) implies approximate-correctness: Suppose $P_{\theta}$ is $\beta$-Self-Proving model with respect to a verifier $V$ that has soundness error $s$ and threshold parameter $\lambda$ for loss function $\ell$. Then by a union bound,
@@
Importantly, as emphasized throughout this paper, soundness of $V$ implies that for all inputs $x$, any output $y$ such that $\ell(x,y)>\lambda$ is rejected with high probability ($1-s$).
Appendix B Theoretical analyses for Section 4
In this section we provide a formal description and analysis of Transcript Learning (TL, Section 4.1) and Reinforcement Learning from Verifier Feedback (RLVF, Section 4.2). In Section B.1 we prove a convergence theorem for TL under convexity and Lipschitzness assumptions. Obtaining an analogous result for RLVF is more challenging; in lieu of a full analysis, we provide a lemma showing that the gradients estimated in the algorithm approximate the Verifiability of the model in Section B.2.
Specification of the learning model.
We must first fully specify the theoretical framework in which our results reside. Continuing from Section 3, we define a learner as an algorithm $\Lambda$ with access to a family of autoregressive models $\{P_{\theta}\}_{\theta}$ and samples from the input distribution $x\sim\mu$. In our setting of Self-Proving models (and in consistence with the Interactive Proofs literature), we give the learner the full specification of the verifier $V$. More formally,
Definition B.1 (Self-Proving model learner).
  • A family of autoregressive models $\{P_{\theta}\}_{\theta\in\mathbb{R}^{d}}$ where $d\in\mathbb{N}$ is the number of parameters in the family. Recall (Section 4) that for each $\theta$ and $z\in\Sigma^{\ast}$, the random variable $P_{\theta}(z)$ is determined by the logits $\log p_{\theta}(z)\in\mathbb{R}^{|\Sigma|}$. For any $z\in\Sigma^{\ast}$ and $\sigma\in\Sigma$, the learner $\Lambda$ can compute the gradient of the $\sigma^{\text{th}}$ logit, that is, $\nabla_{\theta}\log\Pr_{\sigma^{\prime}\sim p_{\theta}(z)}[\sigma=\sigma^{\prime}]$. In particular, $\log\Pr_{\sigma^{\prime}\sim p_{\theta}(z)}[\sigma=\sigma^{\prime}]$ is always differentiable in $\theta$.
@@ Appendix B Theoretical analyses for Section 4
Throughout this section, we will refer to the transcript of an interaction between a verifier and a prover (see Figure 1). We will denote this transcript by $\pi=(y,q_{1},a_{1},\dots,q_{R},a_{R})$, and for any index $s\in[|\pi|]$ we will write $\pi_{<s}\in\Sigma^{s-1}$ to denote the $s$-long prefix of $\pi$.
B.1 Transcript Learning
Recall that Transcript Learning requires access to an honest transcript generator. Before we can formally define this object, it will be useful to define a query generator for a verifier $V$.
Definition B.2 (Query generator).
@@ B.1 Transcript Learning
Fix a verifier $V$ in a proof system with $R\in\mathbb{N}$ rounds, where the verifier issues queries of length $L_{q}=|q_{i}|$ and the prover (model) responses with answers of length $L_{a}=|a_{i}|$. [^8] We can assume that queries (resp. answers) all have the same length by padding shorter ones. The query generator $V_{q}$ corresponding to $V$ takes as input a partial interaction and samples from the distribution over next queries by $V$. Formally, for any $r\leq R$, given input $x$, output $y$, and partial interaction $(q_{i},a_{i})_{i=1}^{r}$, $V_{q}(x,y,q_{1},a_{1},\dots,q_{r},a_{r})$ is a random variable over $\Sigma^{L_{q}}$. [^9] For completeness’ sake, we can say that when prompted with any sequence $z$ that does not encode an interaction, $V_{q}(z)$ is fully supported on a dummy sequence $\bot\cdots\bot\in\Sigma^{L_{q}}$.
A transcript generator is a random variable over transcripts that faithfully represents the interaction of the verifier with some prover for a given input. An honest transcript generator is one who is fully supported on transcripts accepted by the verifier. We denote accepting transcripts by $\pi^{*}=(y^{*},q^{*}_{1},a^{*}_{1},\dots,q^{*}_{R},a^{*}_{R})$.
Definition B.3 (Transcript generator).
Fix a verifier $V$ in a proof system of $R\in\mathbb{N}$ rounds. A transcript generator $\mathcal{T}_{V}$ for $V$ is a randomized mapping from inputs $x\in\Sigma^{\ast}$ to transcripts $\pi=(y,q_{1},a_{1},\dots,q_{R},a_{R})\in\Sigma^{\ast}$. For any input $x$, $\mathcal{T}_{V}(x)$ satisfies that for each $r\leq R$, the marginal of $\mathcal{T}_{V}(x)$ on the $r^{\text{th}}$ query $q_{r}$ agrees with the corresponding marginal of the query generator $(V_{q})_{r}$.
A transcript generator $\mathcal{T}^{\ast}_{V}\coloneqq\mathcal{T}_{V}$ is honest if it is fully supported on transcripts $\pi^{\ast}$ for which the verifier accepts.
@@
Notice that for any verifier $V$, there is a one-to-one correspondence between transcript generators and (possibly randomized) provers. We intentionally chose not to specify a prover in Definition B.3 to emphasize that transcripts can be “collected” independently of the honest prover (see completeness in Definition 3.2), and in fact can be collected “in advance” prior to learning (see Figure 4). As long as the generator is fully supported on honest transcripts, it can be used for Transcript Learning (Algorithm 1 described next).
@@ 9      Update
Figure 4: Transcript Learning, visualized. To understand Algorithm 1, consider the above visualization. In Phase 1, $N$ honest transcripts are collected by letting an Honest Prover interact with the Verification Algorithm; these will be the samples from the honest transcript generator $\mathcal{T}^{\ast}_{V}(x)$. Phase 2 describes the execution of Algorithm 1 itself: For each honest transcript $\pi^{\ast}$ (lines 2-3), and for each prefix $\pi_{s}$ of this transcript (lines 4-6), the $\alpha_{s}(\theta_{i})$ and $\vec{d_{s}}(\theta_{i})$ are computed via forwards and backwards passes, respectively (line 7). After iterating through all prefixes, the parameters $\theta_{i}$ are updated (line 8).
Convergence of TL is proven by a reduction to Stochastic Gradient Descent (SGD). Essentially, we are tasked with proving that TL estimates a surrogate of the Verifiability-gradient of its model $P_{\theta}$. More precisely, TL estimates the gradient of a function that bounds the Verifiability from below. Maximizing this function therefore maximizes the Verifiability.
The lower-bounding function is the agreement of the answers generated by $P_{\theta}$ with the answers provided by the honest transcript generator $\mathcal{T}^{\ast}_{V}$. More formally, we let $\mathcal{T}^{\theta}_{V}$ denote the transcript generator induced by the model $P_{\theta}$ when interacting with $V$: for each $x$, $\mathcal{T}^{\theta}_{V}(x)$ is the distribution over transcripts of interactions between $V$ and $P_{\theta}$ on input $x$. We stress that $\pi^{*}\sim\mathcal{T}^{\ast}_{V}(x)$ and $\pi\sim\mathcal{T}^{\theta}_{V}(x)$ are transcripts produced when interacting with the same verifier queries; we can think of the verifier as simultaneously interacting with the honest prover and with the model $P_{\theta}$. [^10] The way it is presented in the algorithm (and implemented in the experiments), first the verifier is called by $\mathcal{T}^{\ast}_{V}$ and outputs queries $(q^{*}_{1},\dots q^{*}_{R})$, and then the model is prompted with the verifier queries one a time. This maintains soundness, since a proof system is sound as long as the prover does not know the verifier’s queries in advance. In what follows, we use $\pi^{*}\sim\mathcal{T}^{\ast}_{V}(x)$ and $\pi\sim\mathcal{T}^{\theta}_{V}(x)$ to denote two transcripts that share the same queries. That is, taking $\pi^{\ast}=(y^{\ast},q_{1}^{\ast},a_{1}^{\ast},\dots,q_{R}^{\ast},a_{R}^{\ast})$ to denote an accepting transcript sampled from $\mathcal{T}^{\ast}_{V}(x)$, and $\pi=(y,q^{*}_{1},a_{1},\dots,q^{*}_{R},a_{R})$ to denote a random transcript sampled from $\mathcal{T}_{V}^{\theta}(x)$, we say that $\pi$ and $\pi^{*}$ agree if they agree on the prover answers, namely if:
$(y,a_{1},\dots,a_{R})=(y^{\ast},a_{1}^{\ast},\dots,a_{R}^{\ast}).$
This definition implicitly uses the independence of the verifier and model’s randomness. We first prove that TL correctly estimates the gradient of $A(\theta)$ in its update step.
Lemma B.4 (TL gradient estimation).
@@
Fix an input distribution $\mu$ over $\Sigma^{\ast}$ and a verifier $V$ with round complexity $R$ and answer length $L_{a}$. Fix an honest transcript generator $\mathcal{T}_{V}^{\ast}$. Let $\theta$ be the parameters of a model $P_{\theta}$ and let
$A(\theta)\coloneqq\Pr_{\begin{subarray}{c}x\sim\mu\\ \pi^{*}\sim\mathcal{T}^{\ast}_{V}(x)\\ \pi\sim\mathcal{T}^{\theta}_{V}(x)\end{subarray}}\left[\pi=\pi^{*}\right].$
Then,
@@
$\nabla A(\theta)=\operatorname*{\mathbb{E}}_{\begin{subarray}{c}x\sim\mu\\ \pi^{\ast}\sim\mathcal{T}^{\ast}_{V}\end{subarray}}\left[\prod_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\alpha_{s}(\theta)\cdot\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\vec{d}_{s}(\theta)\right]$
where $S(r)$, $\alpha_{s}(\theta)$ and $\vec{d}_{s}(\theta)$ are as defined in Algorithm 1.
@@
Note that Lemma B.4 is true for any model $P_{\theta}$. Moreover, the random vector over which the expectation is taken (in the right hand side) is precisely the direction of the update performed in Algorithm 1. We now prove Lemma B.4, from which we derive Theorem 4.1.
Proof.
Throughout this proof, expectations and probabilities will be over the same distributions as in the lemma statement. First, we use the law of total probability together with the autoregressive property of $P_{\theta}$ (Section 4) to switch from probabilities on transcripts, to products of next-token probabilities. Formally, consider a fixed input $x$, an honest transcript $\pi^{\ast}=(y^{\ast},q_{1}^{\ast},a_{1}^{\ast},\dots,q_{R}^{\ast},a_{R}^{\ast})$, and denote a random transcript sampled from $\mathcal{T}_{V}^{\theta}(x)$ when using the same verifier queries by $\pi=(y,q_{1}^{\ast},a_{1},\dots,q_{R}^{\ast},a_{R})$. For any $r\in[R]$ denote the random variable $\mathcal{T}^{\theta,<r}_{V}\coloneqq\mathcal{T}^{\theta}_{V}(yq^{\ast}_{1}a_{1}\cdots a_{r-1}q^{\ast}_{r})$. Then,
where, as noted above, Equation 3 uses the independence of the verifier and model’s randomness, Equation 4 uses the autoregressive property of $P_{\theta}$ (Definition B.1), and Equation 5 is by definition of $\alpha_{s}$ and of $a_{0}$. Next, a basic calculus identity gives
This implicitly assumes that $\Pr_{\pi}\left[\pi=\pi^{*}\right]$ is differentiable in $\theta$; indeed, this follows from Definition B.1, where the logits of the model were assumed to by differentiable. Let us focus on the rightmost factor. By Equation 5,
where the last equality is by definition of $\vec{d}_{s}(\theta)$. Combining Equation 5 and Equation 6 gives
By the law of total probability and the linearity of the gradient,
which concludes the proof. ∎
We are now ready to prove Theorem 4.1. We restate it below in full formality.
Theorem B.5 (Theorem 4.1, formal).
Fix a verifier $V$, an input distribution $\mu$, an autoregressive model family $\{P_{\theta}\}_{\theta\in\mathbb{R}^{d}}$, and a norm |||||$\cdot$||| | ⋅ | | on $\mathbb{R}^{d}$. Fix an honest transcript generator $\mathcal{T}^{\ast}_{V}$, and assume that the agreement function
$A(\theta)\coloneqq\Pr_{\begin{subarray}{c}x\sim\mu\\ \pi^{*}\sim\mathcal{T}^{\ast}_{V}(x)\\ \pi\sim\mathcal{T}^{\theta}_{V}(x)\end{subarray}}\left[\pi=\pi^{*}\right]$
is concave in $\theta$, where the verifier queries are the same in $\pi^{*}$ and $\pi$. For any $\varepsilon>0$, let $B_{\mathrm{Norm}}$, $B_{\mathrm{Lip}}$ and $C$ be upper-bounds such that the following conditions hold.
  • There exists $\theta^{\ast}\in\mathbb{R}^{d}$ with $||\theta^{\ast}||<B_{\mathrm{Norm}}$ such that $A(\theta^{\ast})\geq 1-\varepsilon/2$.
  • For all $\theta$, the logits of $P_{\theta}$ are $B_{\mathrm{Lip}}$-Lipschitz in $\theta$. That is,
    $\sup_{\begin{subarray}{c}\theta\in\mathbb{R}^{d}\\ z\in\Sigma^{\ast}\end{subarray}}||\nabla_{\theta}\log p_{\theta}(z)||\leq B_{\mathrm{Lip}}.$
Denote by $\bar{\theta}$ the output of TL running for number of iterations $N$ where
$N\geq 4\cdot C^{2}\cdot\frac{B_{\mathrm{Norm}}^{2}\cdot B_{\mathrm{Lip}}^{2}}{\varepsilon^{2}}$
and learning rate $\lambda=B_{\mathrm{Norm}}/CB_{\mathrm{Lip}}\sqrt{N}$. Then the expected Verifiability (over the randomness of the samples collected by TL) of $\bar{\theta}$ is at least $1-\varepsilon$. That is,
$\operatorname*{\mathbb{E}}_{\bar{\theta}}[\mathrm{ver}_{V,\mu}(\bar{\theta)}]\geq 1-\varepsilon.$
Proof.
Our strategy is to cast TL as Stochastic Gradient Ascent and apply C.2. Let $\varepsilon$, $B_{\mathrm{Norm}}$, $B_{\mathrm{Lip}}$ and $C$ as in the theorem statement be given. Let $\theta^{\ast}$ be such that $A(\theta^{\ast})\geq 1-\varepsilon/2$ and $||\theta^{\ast}||\leq B_{\mathrm{Norm}}$.
First, notice that
$\operatorname*{\mathbb{E}}_{\bar{\theta}}\left[\mathrm{ver}_{V,\mu}(\bar{\theta})\right]\geq\operatorname*{\mathbb{E}}_{\bar{\theta}}[A(\bar{\theta})],$
This is because, for any $x$ and model $P_{\theta}$, whenever the transcript generated by $\mathcal{T}^{\theta}(x)$ agrees with $\pi^{\ast}$, then the verifier accepts (because $\pi^{\ast}$ is honest). Therefore, to prove the theorem it suffices to show that
Following the notation in Algorithm 1, in every iteration $i\in[N]$ the norm of the update step is
where the inequality is because $\alpha_{s}(\theta_{i})$ are probabilities, so $\leq 1$. Continuing, we have
To conclude, Lemma B.4 shows that TL samples from a gradient estimator for $A(\theta)$, while the above equation shows that the gradient is upper-bounded by $C\cdot B_{\mathrm{Lip}}$. We can therefore apply C.2 to obtain
where the inequality is by definition of $\theta^{\ast}$.
B.2 Reinforcement Learning from Verifier Feedback
Our second learning method, Reinforcement Learning from Verifier Feedback (RLVF, Algorithm 2), does not require access to an honest transcript generator. Instead, the learner generates transcripts herself by emulating the interaction of the verifier with the current Self-Proving model $P_{\theta}$. When an accepting transcript is generated, the learner updates the parameters $\theta$ towards generating such transcript.

### v4

@@ 3.3.1 The General Definition
Definition 3.7 (Definition˜3.2, generalized).
Fix a soundness error $s\in(0,1)$, a threshold $\lambda\in[0,1)$, a finite set of tokens $\Sigma$, and a loss function $\ell\colon\Sigma^{\ast}\times\Sigma^{\ast}\to[0,1]$. A verifier $V$ for $\ell$ with threshold $\lambda$ is a probabilistic polynomial-time algorithm that is given explicit inputs $x,y\in\Sigma^{\ast}$ and black-box (oracle) query access to a prover $P$. It interacts with $P$ over $R$ rounds (see Figure˜1) and outputs a decision $\left<V,P\right>(x,y)\in\{\mathrm{reject},\mathrm{accept}\}$. Verifier $V$ satisfies the following two guarantees:
@@
Indeed, for a given ground-truth function $F^{\ast}\colon\Sigma^{\ast}\to\Sigma^{\ast}$, Definition˜3.2 can be recovered by choosing the 0-1 loss
$\ell_{F^{\ast}}(x,y)\coloneqq\begin{cases}1\quad\mathrm{if}\ x\neq F^{\ast}(y)\\ 0\quad\mathrm{else}\end{cases}$
@@
Remark 3.8 (Connection to Interactive Proofs of Proximity).
Definition˜3.7 can be seen as a slight generalization of (perfect completeness) Interactive Proofs of Proximity (IPPs, RothblumVW13). An IPP for a relation $L\subseteq\Sigma^{\ast}\times\Sigma^{\ast}$ with proximity parameter $\lambda$ is obtained by instantiating Definition˜3.7 with the loss function $\ell_{\mathrm{Hamming}}$ defined by
@@
that is, $\ell_{\mathrm{Hamming}}(x,y)$ is the fraction of tokens in $y$ that must be changed to obtain an output $y^{\ast}$ with $(x,y^{\ast})\in L$. However, the motivation of RothblumVW13 was studying sublinear time verification, whereas ours is to relax the requirements of traditional Interactive Proofs towards meeting common desiderata in machine learning.
@@
Definition 3.9 (Definition˜3.4, generalized).
Fix a loss function $\ell\colon\Sigma^{\ast}\times\Sigma^{\ast}\to[0,1]$, a verifier $V$ for $\ell$ with threshold $\lambda\in[0,1)$ as in Definition˜3.7, and a distribution $\mu$ over inputs $\Sigma^{\ast}$. The Verifiability of a model $P_{\theta}\coloneq\Sigma^{\ast}\to\Sigma^{\ast}$ is defined as
@@
Analogously to Remark˜3.5, we observe that Verifiability as per Definition˜3.9 implies approximate-correctness: Suppose $P_{\theta}$ is $\beta$-Self-Proving model with respect to a verifier $V$ that has soundness error $s$ and threshold parameter $\lambda$ for loss function $\ell$. Then, by a union bound,
@@
Importantly, as emphasized throughout this paper, soundness of $V$ implies that for all inputs $x$, any output $y$ such that $\ell(x,y)>\lambda$ is rejected with high probability (1s1-s).
4 Algorithms for Learning Self-Proving Models
We focus on differentiable autoregressive models, and assume that the learner has access to input samples $x\sim\mu$ and correct outputs $F^{\ast}(x)$, as well as the verifier’s specification (code). Additionally, the learner can emulate the verifier, as the latter is computationally efficient (Remark˜3.3).
Importantly, we may not assume that the verifier $V$ is differentiable—it is an arbitrary (efficient oracle) Turing machine—and so we cannot directly compute gradients of its decision with respect to model parameters. The challenge is to align the model with a verifier. Algorithms˜1 and 2 address this challenge by (essentially) computing unbiased estimators for the Verifiability $\mathrm{ver}_{V}(\theta)$ or a surrogate (lower-bound) thereof. We formally prove these properties in Lemmas˜4.2 and 4.3.
Our approach is inspired by Reinforcement Learning from Human Feedback (ChristianoLBMLA17), a method for aligning models with human preferences, which has recently been used to align sequence-to-sequence models (Ouyang0JAWMZASR22). However, there are two important differences between humans and algorithmic verifiers: (1) Verifiers are efficient algorithms which may be emulated by the learner. This is unlike humans, whose preferences are costly to obtain. On the other hand, (2) verifiers make a single-bit decision at the end of an interaction, but cannot guide the prover (model) in intermediate rounds. In RL terms, this is known as the exploration problem for sparse reward signals (e.g. LadoszWKO22).
@@ 4 Algorithms for Learning Self-Proving Models
The full specification of the learning model can be found in Section˜A.1. We will refer to the transcript of an interaction between a verifier and a prover (see Figure˜1), denoted by $\pi=(y,q_{1},a_{1},\dots,q_{R},a_{R})$. Let $\pi_{<s}\in\Sigma^{s-1}$ denote the $s$-token long prefix of $\pi$.
4.1 Transcript Learning
We first present an algorithm for learning Self-Proving models which relies on access to a distribution of accepting transcripts. We focus on the algorithm first, and then discuss how the learner may obtain accepting transcripts in Section˜4.1.1. The idea is to learn a model not just of $x\mapsto y^{\ast}$ for a correct output $y^{\ast}$, but of $x\mapsto y^{\ast}\pi^{\ast}$, where $\pi^{\ast}$ is a transcript of an interaction in which the verifier accepted. Formally, Transcript Learning assumes access to a transcript generator—a random variable over transcripts that faithfully represents the interaction of the verifier with some prover for a given input. An honest transcript generator is one which is fully supported on transcripts accepted by the verifier. These are defined next.
Definition 4.1 (Transcript generator).
@@ 4.1 Transcript Learning
Fix a verifier $V$ in a proof system of $R\in\mathbb{N}$ rounds. A transcript generator $\mathcal{T}_{V}$ for $V$ is a randomized mapping from inputs $x\in\Sigma^{\ast}$ to transcripts $\pi=(y,q_{1},a_{1},\dots,q_{R},a_{R})\in\Sigma^{\ast}$. For any input $x$, $\mathcal{T}_{V}(x)$ satisfies that for each $r\leq R$, the marginal of $\mathcal{T}_{V}(x)$ on the $r^{\text{th}}$ query $q_{r}$ agrees with the corresponding marginal of the query generator $(V_{q})_{r}$. [^5] A query generator $V_{q}$ corresponding to $V$ takes as input a partial interaction and samples from the distribution over next queries by $V$. Formally, for any $r\leq R$, given input $x$, output $y$, and partial interaction $(q_{i},a_{i})_{i=1}^{r}$, $V_{q}(x,y,q_{1},a_{1},\dots,q_{r},a_{r})$ is a random variable over $\Sigma^{L_{q}}$. For completeness’ sake, we can say that when prompted with any sequence $z$ that does not encode an interaction, $V_{q}(z)$ is fully supported on a dummy sequence $\bot\cdots\bot\in\Sigma^{L_{q}}$. A transcript generator $\mathcal{T}^{\ast}_{V}\coloneqq\mathcal{T}_{V}$ is honest if it is fully supported on transcripts $\pi^{\ast}$ for which the verifier accepts.
@@
Notice that for any verifier $V$, there is a one-to-one correspondence between transcript generators and (possibly randomized) provers. We intentionally chose not to specify a prover in Definition˜4.1 to emphasize that transcripts can be “collected” independently of the honest prover (see completeness in Definition˜3.2), and in fact can be collected “in advance” prior to learning (see Figure˜2). As long as the generator is fully supported on honest transcripts, it can be used for Transcript Learning as depicted in Algorithm˜1 and Figure˜2.
@@ 9  Update
Figure 2: Transcript Learning, visualized. To understand Algorithm˜1, consider the above visualization. In Phase 1, NN honest transcripts are collected by interacting an honest prover with the Verifier; these serve as samples from the transcript generator $\mathcal{T}^{\ast}_{V}(x)$. Phase 2 runs Algorithm˜1: for each transcript $\pi^{\ast}$ (lines 2–3) and each prefix $\pi_{s}$ (lines 4–6), the values $\alpha_{s}(\theta_{i})$ and $\vec{d}_{s}(\theta_{i})$ are computed via forward and backward passes (line 7), followed by a parameter update (line 8).
TL trains a Self-Proving model by autoregressively optimizing towards generating accepting transcripts. At a high level, it works by repeatedly sampling $x\sim\mu$ and $y^{\ast}\pi^{\ast}\sim\mathcal{T}^{\ast}(x)$, and updating the logits $\log p_{\theta}$ towards agreeing with $y^{\ast}\pi^{\ast}$ via Gradient Ascent. While TL does not directly estimate the Verifiability gradient $\nabla_{\theta}\mathrm{ver}_{V}(\theta)$, we are able to show that it estimates a gradient of a lower-bounding function $A(\theta)\leq\mathrm{ver}_{V}(\theta)$. Therefore, as it ascends the gradient, it optimizes $\mathrm{ver}_{V}(\theta)$ via the surrogate. Formally, the lower-bounding function $A(\theta)$ is the agreement of the transcripts generated by the current model $P_{\theta}$, with the transcripts generated by the honest transcript generator. That is, $A(\theta)\coloneqq\Pr[\pi=\pi^{*}]$ where the probability is over $x\sim\mu$, $\pi_{\theta}\sim\mathcal{T}^{\theta}_{V}(x)$, and $\pi^{*}\sim\mathcal{T}^{\ast}_{V}(x)$.
Lemma 4.2 (TL gradient estimation).
@@
Fix an input distribution $\mu$ over $\Sigma^{\ast}$ and a verifier $V$ with round complexity $R$ and answer length $L_{a}$. Fix an honest transcript generator $\mathcal{T}_{V}^{\ast}$. Let $\theta$ denote the parameters of the model $P_{\theta}$, let $A(\theta)$ be as defined above and let the terms $S(r)$, $\alpha_{s}(\theta)$, and $\vec{d}_{s}(\theta)$ be as defined in Algorithm˜1. Then,
@@
$\nabla A(\theta)=\operatorname*{\mathbb{E}}_{\begin{subarray}{c}x\sim\mu\\ \pi^{\ast}\sim\mathcal{T}^{\ast}_{V}\end{subarray}}\left[\prod_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\alpha_{s}(\theta)\cdot\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\vec{d}_{s}(\theta)\right].$
@@
The proof is deferred to Section˜A.2. Note that Lemma˜4.2 is true for any model $P_{\theta}$. Moreover, the random vector over which the expectation is taken (in the right hand side) is precisely the direction of the update performed in Algorithm˜1. In Section˜5, we will use Lemma˜4.2 to prove convergence bounds for TL under certain conditions.
4.1.1 Access to Accepting Transcripts
As mentioned, Transcript Learning relies on access to accepting transcripts. In this section we discuss how such access can be realized (grounded in the theory of Interactive Proofs).
Doubly-Efficient Interactive Proofs.
When the honest prover $P^{\ast}$ is efficient (polynomial time), the learner (who has the code for $P^{\ast}$ and $V$) can execute $P^{\ast}$ on input $x$ to collect accepting transcripts—asuming no distribution shift at inference time. This setting is formalized by the notion of Doubly-Efficient Interactive Proofs (DEIPs), introduced in the foundational work of  GKR08, who construct DEIPs for all problems computable by log-space uniform families of polynomial-size circuits with polylogarithmic depth. Their protocols ensure that the verifier runs in nearly linear time, while the prover operates in polynomial time—or more generally, in time proportional to the circuit size. Later,  RRR16 showed that any problem computable by a Turing machine running in polynomial time and sublinear space admits a constant-round DEIP. Subsequent theoretical works (e.g., GR18; GoldreichR18) have constructed DEIPs for specific problems, while applied works (e.g., ZhangLWZSXZ21; Thaler) have improved the time and space complexity of such protocols in practice.
Backward Data Generation.
Recent works on “AI for advanced mathematical problems” (charton2021; Lyapunov) propose to reverse the generation process of problem-solution pairs as follows: rather than sampling problems and searching for solutions, one first samples solutions and then constructs the corresponding problems. In our setting, this inspires the following approach. Suppose it is computationally hard in the worst case to start with the input $x$ and generate an accepting transcript $\pi=(y,q_{1},a_{1},\dots,q_{R},a_{R})$ for $x$. Could we instead jointly sample $(x,\pi)$, or first sample a transcript $\pi$ and then extract the input $x$ for which $\pi$ is an accepting transcript? As long as the resulting $x$’s are distributed the same as inference time, this would enable Transcript Learning of such instances.
We explain the idea further with a cryptographic example. Setting up a Diffie–Hellman key-exchange scheme for security parameter $1^{k}$ requires producing as a public parameter a prime $p$ in a factored form, namely $p-1=\Pi{{q_{i}}^{\alpha_{i}}}$ where $q_{i}$ are primes. However, factoring (p1)(p-1) is generally hard. Instead, one could go “backwards:” first generate (p1)(p-1) in factored form by choosing primes $q_{i}$ and exponents $\alpha_{i}$ (kalai2003generating) and then testing if $p=1+\Pi{{q_{i}}^{\alpha_{i}}}$ is prime. By the prime number theorem, $p$ is likely to be prime after a few attempts. In the context of our paper, one may ask a Self-Proving model to produce certified primes, training it on tuples $(x\coloneqq p,\pi\coloneqq(q_{i},\alpha_{i})_{i})$.
4.2 Reinforcement Learning from Verifier Feedback (RLVF)
As mentioned in Section˜4.1, Transcript Learning uses access to an honest transcript generator to estimate gradients of (a lower bound on) the Verifiability of a model $P_{\theta}$. Next we present Reinforcement Learning from Verifier Feedback (RLVF, Algorithm˜2), which estimates this gradient without access to a transcript generator.